OpenStudy (waheguru):

Is is true when one part of a fraction is negative the whole fraction becomes negative

OpenStudy (waheguru):

\[60/-12 = 5? or -5 \]

OpenStudy (unklerhaukus):

\[\frac {-a}b=\frac{a}{-b}=-\frac ab\]

OpenStudy (waheguru):

This is from the equation −3(4r − 8) = −36 I am getting -5 but the real answer is 5 ?

OpenStudy (unklerhaukus):

\[−3(4r − 8) = −36 \] \[4r − 8 = \frac{−36}{-3} \]

OpenStudy (unklerhaukus):

\[\frac {-a}{-b}=\frac{a}{b}\]

OpenStudy (waheguru):

Rather than that step...

OpenStudy (waheguru):

Cant I distrubute the -3 and get -12r+24=-36

OpenStudy (unklerhaukus):

that way will work too

OpenStudy (waheguru):

So how Do I get 5 rather than -5

OpenStudy (unklerhaukus):

using the distributive method the next step would be to get the constants of one side

OpenStudy (waheguru):

yea I have ...

OpenStudy (waheguru):

-12r = 60

OpenStudy (waheguru):

and then ...

OpenStudy (waheguru):

r = 60/-12 r= -5?

OpenStudy (unklerhaukus):

-12r = -60

OpenStudy (waheguru):

-60?

OpenStudy (waheguru):

yea yout correct

OpenStudy (unklerhaukus):

you took 24 away from both sides

OpenStudy (waheguru):

then its r=-60/-12 and ots r=5

OpenStudy (waheguru):

Thanks!

OpenStudy (unklerhaukus):

that is right \[ \color\red\checkmark \]